What is the density of propene gas measured at pressure of 0.5 atm and temperature of 470C?
1
Expert's answer
2012-12-06T09:38:34-0500
470 C = 273 + 470 K = 743 K
first we find concentration of the propene:
p = nkT
n = p/Tk = 0.5* 101.3*103 Pa/ (743 K 1.38*10^(-23)m^2 s^(-2)*K^(-1)) =
49.4*10^(23) molecules/ m^3

now to find density
ho = n* (mass of molecule) = 49.4*10^(23) * (42.08*1.a.m.u) = 49.4*10^(23) * 42.08* 1.660*10^(-27) = 0.345 kg/m^3

density of propene is 0.345 kg/m^3